OpenStudy (anonymous):

Line AB goes through the points A (0, –4) and B (6, 2). Which equation represents line AB? y – 4 = 3x y – 2 = 3(x – 6) y + 4 = x y + 6 = x + 2

OpenStudy (anonymous):

To find the equation of the line, first find the slope through the slope formula. \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] Plug in the points into the equation. \[m=\frac{ 2+4 }{ 6-0 }\] m = 6/6 m = 1 Plug that into the y - y_1 = m(x - x_1) formula. y + 4 = 1(x - 0) y + 4 = x The third choice is correct.